NASA ISS On-Orbit Status 16 April 2011


image All ISS systems continue to function nominally, except those noted previously or below. Saturday - Crew rest day.

Upon wake-up, CDR Kondratyev performed the regular daily check of the aerosol filters at the Russian Elektron O2 (oxygen) generator which Maxim Suraev had installed on 10/19/09 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [Dmitri will inspect the filters again before bedtime tonight, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

FE-1 Samokutyayev completed his 2nd session of the MedOps PZE MO-2 24-hour protocol with ECG (electrocardiograph) & BP (blood pressure) measurements. [After the ECG recording and BP measurements with the Kardiomed system, Alexandr doffed the five-electrode Holter harness that read his dynamic (in motion) heart function from two leads over the past 24 hours, recording data on the "Kardioregistrator 90205" unit. The examination results were then downloaded from the Holter ECG device to the RSE-Med laptop, controlled by the Kardiomed application. Later, the data were downlinked as a compressed .zip-file via OCA.]

FE-3 Garan, FE-5 Nespoli & FE-6 Coleman completed another post-sleep session of the Reaction Self Test (Psychomotor Vigilance Self Test on the ISS) protocol. It was the 3rd for Ron, the 26th for Paolo & Cady. [RST is done twice daily (after wakeup & before bedtime) for 3 days prior to the sleep shift, the day(s) of the sleep shift and 5 days following a sleep shift. The experiment consists of a 5-minute reaction time task that allows crewmembers to monitor the daily effects of fatigue on performance while on ISS. The experiment provides objective feedback on neurobehavioral changes in attention, psychomotor speed, state stability, and impulsivity while on ISS missions, particularly as they relate to changes in circadian rhythms, sleep restrictions, and extended work shifts.]

The crew joined in conducting the regular weekly three-hour task of thorough cleaning of their home, including COL (Columbus Orbital Laboratory) and Kibo JPM (JEM Pressurized Module). ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the uborka house cleaning, CDR, FE-1 & FE-2 completed regular maintenance inspection & cleaning of fan screens in the FGB (TsV2) plus Group E fan grilles (VPkhO, FS5, FS6, VP) and the BMP Harmful Contaminants Removal System grille in the SM (Service Module).

Kondratyev conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and filling EDV-SV, KOV (for Elektron), EDV-ZV & EDV on RP flow regulator.]

FE-5 Nespoli worked on the ATV (Automated Transfer Vehicle) hatch at the SM aft port, securing it on its intended support points using an adjustable bungee assembly from Node-1 stowage. [The hatch has proven difficult to be positioned as intended. New instructions were uplinked to Paolo for an alternate method of securing it.]

FE-3 Garan spent ~50min with FE-6 Coleman in another handover session, their 4th, during which Cady familiarized her compatriot with USOS (US Segment) activities.

At ~8:20am, Alex Samokutyayev had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

At ~9:00am EDT, the six crewmembers held the regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP-Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

At ~1:10pm, Dima, Sasha & Andrey engaged in a PAO phone interview via S-band with Ekaterina Beloglazova, Editor of Rossiyskiy Kosmos (Russian Space) Magazine and an old friend of ISS cosmonauts.

Jobs listed for CDR Kondratyev today on the Russian discretionary "time permitting" task list were -
* A ~30-min. session for Russia's EKON Environmental Safety Agency, making observations and taking KPT-3 aerial photography of environmental conditions on Earth using the NIKON D3X camera with the RSK-1 laptop, and
* A ~15-min. photography session for the DZZ-13 "Seiner" ocean observation program, obtaining NIKON D3 photos with Nikkor 80-200 mm lens and the SONY HD video camcorder on oceanic color bloom patterns in the waters of the South-Eastern Pacific, then copying the images to the RSK-1 laptop.

The crew worked out with their regular 2-hr physical exercise on the TVIS treadmill (CDR, FE-1, FE-2), ARED advanced resistive exercise device (CDR, FE-1, FE-3, FE-5, FE-6), T2/COLBERT advanced treadmill (FE-6) and VELO ergometer bike with load trainer (FE-2, FE-3, FE-5).

Weekly Science Update (Expedition Twenty-Seven -- Week 4)

2D NANO Template (JAXA): No report.

3D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA SHIELD (NASA/ASI): No report.

AMS-02 (Alpha Magnetic Spectrometer): AMS Laptop: Issues with 2nd hard drive resolved from the ground 4/8. The on-orbit procedure will be changed to prevent reoccurrence. Daily commanding and weekly downlinks progressing nicely.

APEX (Advanced Plant Experiments on Orbit) -Cambium: No report.

APEX-TAGES (Transgenic Arabidopsis Gene Expression System): No report.

Asian Seed 2010 (JAXA): Photo session is on the task list.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): "With the present photos and a LOT of image processing we are able to see that the BCAT-5, sample 9 images indicate colloidal crystals forming after about 12 - 15 days. We believe that photos taken at another angle will allow us to convince ourselves and others that we are clearly seeing the effects of 'nanodirt' on colloidal crystal formation."

BIOLAB (ESA): "Dear Cady, we are happy that the BIOLAB Microscope and Spectrophotometer are safely packed and ready for return on ULF6. We really appreciate all your valuable comments on the procedure; these will certainly make new procedures more accurate. We are looking forward to the pictures you took. Thanks for all your work on BIOLAB and for finishing the packing."

BIORHYTHMS (JAXA, Biological Rhythms): No report.

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: No report.

BXF-Facility (Boiling eXperiment Facility, NASA): The dissolved gas content was measured at 712 ppm at the start of operations this week. The written requirement is <100 PPM, although there is still some good publishable data. The flight experiment measures at an accuracy of +/- 130 ppm. After about 12 hours of operations, sensors in the BXF system began generating unexpected readings, and the BXF system placed itself in a safe-state condition. Monitoring was able to continue. Initial troubleshooting results were disappointing, and MSG removed power from the BXF. The ops team will continue to investigate, and determine a course of action.

BXF-MABE (Microheater Array Boiling Experiment, NASA): 55 MABE tests were conducted using the 7.0 mm microheater array prior to the cessation of activities. We accomplished 75% of our science, even though 35% of our test points were completed.

BXF-NPBX (Pool Boiling Experiment, NASA): No activities were planned.

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: No report.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): No report.

CERISE (JAXA): No report.

CCF (Capillary Channel Flow): No report.

CFS-A (Colored Fungi in Space-A, ESA): No report.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack), MDCA/Flex: No report.

Commercial (Inc 23&24, JAXA): No report.

Commercial (Inc 25 & 26, JAXA): Sample launch & return by ULF6.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): No report.

CsPins (JAXA): Will follow Nanoskeleton 2.

CubeLab: No report.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DECLIC-ALI (Device for the Study of Critical Liquids & Crystallization-ALICE-like, CNES/NASA): No report.

DomeGene (JAXA): Complete.

DOSIS (Dose Distribution Inside ISS, ESA): Acquiring science data with DOSTEL-2 detector.

EarthKAM (Earth Knowledge Acquired by Middle School Students): No report.

EDR (European Drawer Rack, ESA): No report.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EMCS (European Modular Cultivation System): No report.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): Activated in support of CARD.

EPO (Educational Payload Operations, NASA) (Hobbies): No report.

EPO (Educational Payload Operations, NASA) (Kids in Space): No report.

EPO LES-2 (ESA): No report.

EPO GREENHOUSE (ESA): No report.

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): No report.

EPO Dewey's Forest (JAXA): Closed out on 3/15.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): No report.

EPO-5 SpaceBottle (Message in a Bottle, JAXA): No report.

EPO Moon Score (JAXA): No report.

EPO Try Zero-G (JAXA): No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Paper Craft (Origami, JAXA): No report.

EPO Poem (JAXA): No report.

EPO Spiral Top (JAXA): No report.

ERB-2 (Erasmus Recording Binocular, ESA): [ERB-2 aims are to develop narrated video material for various PR & educational products & events, including a 3D interior station view.] "Paolo, unfortunately the last downlink on 4/14 was not successful: the team could downlink 1 file; the file transfer between ERB-2 and the EDR VMU slowed down for the second file and had to be stopped. Additional attempts to start the transfer of another file didn't solve the issue. So the cable connection should be fine. The anomaly occurred before and being investigated."

ETD (Eye Tracking Device): Completed.

FACET-2 (JAXA): No report.

FERULATE (JAXA): No report.

FIR/LMM/CVB (Fluids Integrated Rack / Light Microscopy Module / Constrained Vapor Bubble): No report.

Fish Scales (JAXA): Completed on FD7/ULF-4 and returned on STS-132.

FOAM STABILITY (ESA): No report.

FOCUS: No report.

FSL (Fluid Science Laboratory, ESA): Active for GEOFLOW-II experiment. Refer to GEOFLOW-II.

FWED (Flywheel Exercise Device, ESA): No report.

GENARA-A (Gravity Regulated Genes in Arabidopsis A/ESA): No report.

GEOFLOW (ESA): For each of the first executed runs, set-points with high delta_T were skipped by the automatic script due to a PID control time-out on one of the parameters. A second anomaly affecting the science execution is the occurrence of a FSL TEC power board trip. Engineering teams concluded that the hardware can tolerate the trips. Both anomalies are under investigation by the engineering teams. In anticipation of a work-around, the engineering teams gave the go-ahead to execute runs which have a lower risk of tripping and which are restricted to a maximum delta_T which is lower than the desired scientific set-point. Those runs are the so-called "rotation runs" and have the lowest priority for science. Depending on the rotation speed of the experimental fluid cavity, the runs are classified as low (0.008 Hz - 0.16 Hz), medium (0.2 Hz - 1.0 Hz) or high (1.2 Hz - 2.0 Hz) rotation runs. On 4/11, a first low rotation run was started and successfully executed, followed by a successful medium rotation run on 4/12. During the last 2 rotation set-points of the high rotation run on 4/12, the tray did not rotate anymore and the run was aborted. Engineering teams reported that a possible root cause is slippage of the belt for the rotating tray belt, due to low tension on the belt. On 4/13 the engineering teams decided that all rotation runs up to 1.6 Hz are allowed to be executed. Together with the 2 previous anomalies, this further restricts the options parameters ranges (Delta T, rotation speed) to perform science runs for the moment.

HAIR (JAXA): No report.

HDTV System (JAXA): Was delivered by HTV1.

Hicari (JAXA): Prior to Hicari, Kobairo checkout is progressing.

Holter ECG (JAXA): No report.

HQPC (JAXA): Was delivered by 34P.

HREP (HICO/Hyperspectral Imager for the Coastal Ocean & RAIDS/Remote Atmospheric & Ionospheric Detection System/JAXA): HICO has taken 2917 images to date. The most recent HICO images taken include the coast of Japan near the Fukushima Dai-ichi plant, Panama City, FL and the Oaxaca Valley in Mexico. RAIDS is collecting secondary science including nighttime atmospheric disk photometry, spectra and temperatures. Extreme Ultra Violet airglow spectroscopy and optical contamination studies are also being performed.

HydroTropi (Hydrotropism & Auxin-Inducible Gene Expression in Roots Grown under Microgravity Conditions/JAXA): No report.

ICE CRYSTAL (JAXA): Complete.

ICV (Integrated Cardiovascular): No report.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

INTEGRATED IMMUNE: No report.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.

IRIS (Image Reversal in Space, CSA): No report.

ISS Amateur/Ham Radio: No report.

IV Gen (Intravenous Fluids Generation): No report.

KID/KUBIK6: No report.

Kids in Micro-G: No report.

KUBIK 3 (ESA): No report.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.

Marangoni Exp (JAXA): 24th run was completed on 12/22.

Marangoni DSD - Dynamic Surf (JAXA): Payload name was change from Marangoni DSD to Dynamic Surf.

Marangoni UVP (JAXA): We are planning to return MI core and Cassette on ULF7 or SpX1 for troubleshooting.

Matryoshka-2 (RSA): No report.

MAXI (Monitor of All-sky X-ray Image, JAXA): No report.

MDCA/Flex: No report.

MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): No report.

Microbe-2 (JAXA): Sample return by ULF6.

Micro-G Clay (JAXA EPO): Complete.

MISSE-7 (Materials ISS Experiment): MISSE-7 is operating nominally. PEC B is powered off, and awaiting retrieval during STS-134. The CMOS Imager Experiment is testing the radiation effects on a CMOS based imager chip. An image will be downloaded over the next two weeks to examine in detail the nature of radiation degradation. The SpaceCube experiment has the ability to be reconfigured from the ground with updated code and the process of uploading new radiation hardening by software experiment algorithms is in progress. About 40% of the file uploads are complete and the remaining uploads will be spread out over at least another month.

MMA (JAXA/Microgravity Measurement Apparatus): No report.

MPAC/SEED (JAXA): Completed on 19A FD4.

MSG-SAME (Microgravity Science Glovebox-Smoke Aerosol Measurement Experiment): No report.

MSL (Materials Science Laboratory, ESA): Last Sample Cartridge Assembly (SCA) returned with STS-133 (ULF5).

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC-1 "Pirs".

MULTIGEN-1: Completed.

MYCO 3 (JAXA): No report.

MyoLab (JAXA): Completed on 4/20.

NANOSKELETON (Production of High Performance Nanomaterials in Microgravity, JAXA): Experiment Exp2 completed on 4/8.

NEURORAD (JAXA): No report.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration):

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY/ProK: "Ron, you've completed your first of five Nutrition sessions, with all activities successfully accomplished well within the experiment constraints. Paolo, the Pro K team would like to thank you for your patience while the ground team worked the food container issues. All your extra efforts to retain your planned menu was much appreciated by the PI, who reported you did a great job with your selection!"

PADIAC (Pathway Different Activators, ESA): No report.

PADLES (JAXA, Area PADLES 3/4; Passive Area Dosimeter for Lifescience Experiment in Space): No report.

PASSAGES (JAXA): No report.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.

PCG (JAXA, Protein Crystal Growth): No report.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): See PCG.

PMDIS (Perceptual Motor Deficits in Space): Complete.

POLCA/GRAVIGEN (ESA): Complete.

Pro K: No report.

RadGene & LOH (JAXA): Complete.

RadSilk (JAXA): No report.

RST/Reaction Self Test (Psychomotor Vigilance Self Test on the ISS): "Paolo and Cady, thank you for completing the Reaction Self Test sleep shift sequence. The data from this will be very useful."

RYUTAI Rack (JAXA): No report.

SAIBO Rack (JAXA): No report.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): No report.

SAMPLE: Complete.

SCOF (Solution Crystallization Observation Facility, JAXA): No report.

SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): Continuous operation from Inc 19&20.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLAMMD (Space Linear Acceleration Mass Measurement Device): No report.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): No report.

SMILES (JAXA): Recooling mode #13 continues.

SODI/IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in Liquids, ESA): No report.

SODI/COLLOID (Selectable Optical Diagnostics Instrument/Colloid): COLLOID flash-disks returned with STS-133 (ULF5).

SOLAR (Solar Monitoring Observatory): Out of Sun Visibility Window.
The SolACES team reported that they observed a significantly higher degradation of the spectrometers than before. The suspicion is that external venting particles reached the optics. They requested to raise the temperature of the instrument around +42degC to +45degC for the current dark period in the hope this strategy would remove the particles. SolACES temperature was changed on 4/7. Later on, the science team reconsidered their strategy because they are concerned that the efficiency of the 2 affected spectrophotometers might improve, but this continuous heating might have a negative impact on the third spectrophotometer (this spectrophotometer showing the lowest degradation for the moment). To preserve the functionality of the third spectrometer, the temperature of the SolACES instrument was set back to the standard value on 4/13. The effect of the heating can only be evaluated at the start of the next Sun Visibility Window (currently planned on 4/29). The root source of the particles is not yet known and it is important to identify any venting, thrusters, vehicle traffic or any other contamination events from Columbus and/or ISS that occurred from February to March 2011. This investigation is on-going.

SOLO (Sodium Loading in Microgravity): No report.

Space-DRUMS (Space Dynamically Responding Ultrasonic Matrix System): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.

SPHINX (SPaceflight of Huvec: an Integrated eXperiment, ESA): No report.

SPICE (Smoke Point In Co-flow Experiment): No report.

SPINAL (Spinal Elongation): No report.

SWAB (Characterization of Microorganisms & Allergens in Spacecraft): No report.

TASTE IN SPACE (ESA): No report.

THERMOLAB (ESA): "Cady, great job on your 4th session. We would like to quote the words of the science team from ThermoLab side: "The data are wonderful. Please keep going ;-)"

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

TRIPLELUX-B (ESA): No report.

ULTRASOUND: Planned.

VASCULAR (CSA): No report.

VCAM (Vehicle Cabin Atmosphere Module, NASA): No report.

VESSEL ID System (ESA): Acquiring science data with NorAIS receiver.

VESSEL IMAGING (ESA): No sessions possible due to Ultrasound anomaly.

VO2max (NASA): No report.

VLE (Video Lessons ESA): No report.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels; ESA): No report.

YEAST B (ESA): No report.

CEO (Crew Earth Observation): For Increment 27 through 4/13, 5,735 ISS CEO frames were received on the ground for review and cataloging. "We are pleased to report your acquisition of the following target with a time corresponding to one of our daily CEO Target Request list: Aurora Borealis, Alaska, NW Canada - a nice series of views of the rather limited display at the time of your pass - we will continue to ask for views of this beautiful phenomena as conditions permit. As a bonus within your Aurora session you also acquired several crisp views of Edmonton, Alberta. Based on your interest and ability demonstrated in this nice imagery we will begin requesting city-at-night views in our future target lists. Your striking, sun glint-enhanced view of agricultural fields along the Shebelle River, Ethiopia was published on NASA/GSFC's Earth Observatory website this past weekend. This semi-arid plateau region of southeastern Ethiopia is a most unlikely place to spot surface water features, but with your alert timing and exposure, a fairly elaborate agricultural irrigation grid near the river comes sharply into view. Very nice shot!

CEO targets uplinked for today were Bangkok, Thailand, at night (looking left. The lights and coastline of this large port city should have stood out), Woollya Cove, Chile (HMS Beagle site. Looking right among small islands south of Tierra del Fuego. Charles Darwin stopped here in 1833. Overlapping frames of the islands were requested), and Tehran, Iran, at night (looking left of track for this capital city at night).

ISS Orbit (as of this morning, 8:42am EDT [= epoch])
Mean altitude - 348.9 km
Apogee height - 351.0 km
Perigee height - 346.8 km
Period -- 91.52 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0003142
Solar Beta Angle -- 61.4 deg (magnitude increasing)
Orbits per 24-hr. day -- 15.73
Mean altitude loss in the last 24 hours -- 127 m
Revolutions since FGB/Zarya launch (Nov. 98) - 71,119

Significant Events Ahead (all dates Eastern Time and subject to change):
--------------Six-crew operations-------------
04/22/11 -- Progress M-09M/41P undock ~7:41am EDT
04/26/11 -- Progress M-09M/41P deorbit ~9:15am
04/27/11 -- Progress M-10M/42P launch
04/29/11 -- Progress M-10M/42P docking (DC-1 nadir)
04/29/11 -- STS-134/Endeavour launch ULF6 (ELC-3, AMS) ~3:47:49pm EDT
05/01/11 -- STS-134/Endeavour docking ~1:31pm
05/11/11 -- STS-134/Endeavour undocking ~6:23am
05/13/11 -- STS-134/Endeavour landing (KSC) ~9:28am
05/16/11 - Soyuz TMA-20/25S undock/landing (End of Increment 27)
--------------Three-crew operations-------------
05/30/11 -- Soyuz TMA-02M/27S launch - M. Fossum (CDR-29)/S. Furukawa/S. Volkov
06/01/11 -- Soyuz TMA-02M/27S docking (MRM1)
--------------Six-crew operations-------------
06/xx/11 -- ATV-2 "Johannes Kepler" undock (SM aft)
06/21/11 -- Progress M-11M/43P launch
06/23/11 -- Progress M-11M/43P docking (SM aft)
06/28/11 -- STS-135/Atlantis launch ULF7 (MPLM) ~3:30pm EDT NET
06/30/11 -- STS-135/Atlantis docking ULF7 (MPLM) NET
07/27/11 - Russian EVA #29
08/29/11 -- Progress M-11M/43P undocking
08/30/11 -- Progress M-12M/44P launch
09/01/11 -- Progress M-12M/44P docking (SM aft)
09/16/11 - Soyuz TMA-21/26S undock/landing (End of Increment 28)
--------------Three-crew operations-------------
09/30/11 -- Soyuz TMA-03M/28S launch - D.Burbank (CDR-30)/A.Shkaplerov/A.Ivanishin
10/02/11 - Soyuz TMA-03M/28S docking (MRM2)
--------------Six-crew operations-------------
10/25/11 -- Progress M-10M/42P undocking
10/26/11 -- Progress M-13M/45P launch
10/28/11 -- Progress M-13M/45P docking (DC-1)
11/16/11 -- Soyuz TMA-02M/27S undock/landing (End of Increment 29)
--------------Three-crew operations-------------
11/30/11 -- Soyuz TMA-04M/29S launch - O.Kononenko (CDR-31)/A.Kuipers/D.Pettit
12/02/11 -- Soyuz TMA-04M/29S docking (MRM1)
--------------Six-crew operations----------------
12/26/11 -- Progress M-13M/45P undock
12/27/11 -- Progress M-14M/46P launch
12/29/11 -- Progress M-14M/46P docking (DC-1)
02/29/12 -- ATV3 launch readiness
03/05/12 -- Progress M-12M/44P undock
03/16/12 -- Soyuz TMA-03M/28S undock/landing (End of Increment 30)
--------------Three-crew operations-------------
03/30/12 -- Soyuz TMA-05M/30S launch - G.Padalka (CDR-32)/J.Acaba/K.Volkov
04/01/12 -- Soyuz TMA-05M/30S docking (MRM2)
--------------Six-crew operations----------------
05/05/12 -- 3R Multipurpose Laboratory Module (MLM) w/ERA - launch on Proton (under review)
05/06/12 -- Progress M-14M/46P undock
05/07/12 -- 3R Multipurpose Laboratory Module (MLM) - docking (under review)
05/16/12 -- Soyuz TMA-04M/29S undock/landing (End of Increment 31)
--------------Three-crew operations-------------
05/29/12 - Soyuz TMA-06M/31S launch - S.Williams (CDR-33)/Y.Malenchenko/A.Hoshide
05/31/12 - Soyuz TMA-06M/31S docking
--------------Six-crew operations----------------
09/18/12 -- Soyuz TMA-05M/30S undock/landing (End of Increment 32)
--------------Three-crew operations-------------
10/02/12 -- Soyuz TMA-07M/32S launch - K.Ford (CDR-34)/O.Novitskiy/E.Tarelkin
10/04/12 - Soyuz TMA-07M/32S docking
--------------Six-crew operations-------------
11/16/12 -- Soyuz TMA-06M/31S undock/landing (End of Increment 33)
--------------Three-crew operations-------------
11/30/12 -- Soyuz TMA-08M/33S launch - C.Hadfield (CDR-35)/T.Mashburn/R.Romanenko
12/02/12 - Soyuz TMA-08M/33S docking
--------------Six-crew operations-------------
03/xx/13 -- Soyuz TMA-07M/32S undock/landing (End of Increment 34)
--------------Three-crew operations-------------
03/xx/13 - Soyuz TMA-09M/34S launch - P.Vinogradov (CDR-36)/C.Cassidy/A.Misurkin
03/xx/13 - Soyuz TMA-09M/34S docking
--------------Six-crew operations-------------
05/xx/13 - Soyuz TMA-08M/33S undock/landing (End of Increment 35)
--------------Three-crew operations-------------
05/xx/13 - Soyuz TMA-10M/35S launch - M.Suraev (CDR-37)/K.Nyberg/L.Parmitano
05/xx/13 - Soyuz TMA-10M/35S docking
--------------Six-crew operations-------------
09/xx/13 - Soyuz TMA-09M/34S undock/landing (End of Increment 36)
--------------Three-crew operations-------------
09/xx/13 - Soyuz TMA-11M/36S launch - M.Hopkins/TBD (CDR-38)/TBD
09/xx/13 - Soyuz TMA-11M/36S docking
--------------Six-crew operations-------------
11/xx/13 - Soyuz TMA-10M/35S undock/landing (End of Increment 37)
--------------Three-crew operations-------------
11/xx/13 - Soyuz TMA-12M/37S launch - K.Wakata (CDR-39)/R.Mastracchio/TBD
11/xx/13 - Soyuz TMA-12M/37S docking
--------------Six-crew operations-------------
03/xx/14 - Soyuz TMA-11M/36S undock/landing (End of Increment 38)
--------------Three-crew operations-------------

Please follow SpaceRef on Twitter and Like us on Facebook.